# Decoding Knowledge Attribution in Mixture-of-Experts: A Framework of Basic-Refinement Collaboration and Efficiency Analysis

### v1

@@ Abstract
The interpretability of Mixture-of-Experts (MoE) models, especially those with heterogeneous designs, remains underexplored. Existing attribution methods for dense models fail to capture dynamic routing-expert interactions in sparse MoE architectures. To address this issue, we propose a cross-level attribution algorithm to analyze sparse MoE architectures (Qwen 1.5-MoE, OLMoE, Mixtral-8x7B) against dense models (Qwen 1.5-7B, Llama-7B, Mixtral-7B). Results show MoE models achieve 37% higher per-layer efficiency via a “mid-activation, late-amplification” pattern: early layers screen experts, while late layers refine knowledge collaboratively. Ablation studies reveal a “basic-refinement” framework—shared experts handle general tasks (entity recognition), while routed experts specialize in domain-specific processing (geographic attributes). Semantic-driven routing is evidenced by strong correlations between attention heads and experts ($r=0.68$), enabling task-aware coordination. Notably, architectural depth dictates robustness: deep Qwen 1.5-MoE mitigates expert failures (e.g., 43% MRR drop in geographic tasks when blocking top-10 experts) through shared expert redundancy, whereas shallow OLMoE suffers severe degradation (76% drop). Task sensitivity further guides design: core-sensitive tasks (geography) require concentrated expertise, while distributed-tolerant tasks (object attributes) leverage broader participation. These insights advance MoE interpretability, offering principles to balance efficiency, specialization, and robustness.
@@ 1 Introduction
We make three contributions:
@@ 2.1 MoE Architecture
In the decoder, the final hidden states $\bm{h}_{i}^{L}$ need to go through the unembedding process to map the high-dimensional representations back to the vocabulary space. Specifically, for each position $i$, the final output token vector $\bm{y}_{i}$ is obtained by multiplying the hidden state $\bm{h}_{i}^{L}$ with the unembedding matrix $\bm{W}_{\text{unembed}}$. At this point, $\bm{y}_{i}$ represents the model’s predicted distribution over the vocabulary at position $i$.
@@ 2.2 Neuron-Level Attribution
We adopt the neuron-level attribution methods from Yu and Ananiadou (2024) for the FFN and attention layers and extend them for MoE models. For attention neurons, the importance score is:
$\mathcal{I}(\bm{v}^{l}_{A})=\log p(x_{i}|\bm{v}^{l}_{A}+\bm{h}^{l-1})-\log p(x_{i}|\bm{h}^{l-1}).$ (7)
@@
where $\bm{v}^{l}$ is the output of the FFN at layer $l$.
@@
where $g_{i,j}^{l}$ is the gating probability for expert $\mathcal{E}_{j}$, and $v^{l}_{\mathcal{E}_{j}}$ is the output from expert $\mathcal{E}_{j}$.
@@ 3.1 Layer Efficiency and Knowledge Concentration in MoE Models
Table 1 summarizes the performance of the four models, revealing that MoE architectures achieve notably higher per-layer gains. For example, Qwen 1.5-MoE attains a layer efficiency of 0.294 using only 24 layers, whereas its dense counterpart, Qwen 1.5-7B, requires 32 layers to reach an efficiency of just 0.213. Despite a slightly lower overall FFN gain, OLMoE—by virtue of its shallower design—achieves the highest layer efficiency at 0.299, demonstrating a marked optimization in computational density.
The data further indicate that MoE models concentrate their knowledge processing at earlier stages. Specifically, Qwen 1.5-MoE and OLMoE reach their peak FFN gains at 87.8% and 78.7% of the network depth, respectively. This “early routing, late refinement” pattern enables these models to perform initial expert filtering in mid-layers and then focus on intensive knowledge refinement in later layers. In contrast, dense models such as Qwen 1.5-7B and Llama-7B display a more gradual accumulation of gains—peaking at 90.2% and 79.2% of the network depth—which, while beneficial for cross-layer integration, tends to blur the functional distinctions among layers and reduce interpretability.
@@
Furthermore, a comparison of attention versus FFN gains underscores divergent processing strategies. Qwen 1.5-MoE, for instance, exhibits a pronounced reliance on specialized expert networks, with an FFN gain of 7.05 contrasted against an attention gain of only 3.58. Dense models, by comparison, distribute gains more evenly (with attention gains of 3.91 and 4.06 for Qwen 1.5-7B and Llama-7B, respectively), complicating the attribution of specific processing roles to individual layers. Notably, OLMoE achieves nearly identical attention and FFN gains (4.78 and 4.79, respectively), reflecting a balanced strategy that integrates both attention mechanisms and expert routing.
@@ 3.2 FFN Gain Evolution: Mid-Stage Activation and Late-Stage Amplification
Figure 3: Cumulative FFN Layers Contribution Curves of the six models.
@@ 4.3 Causality Analysis
We selectively suppressed the activity of head $H^{19}_{12}$ in Qwen during inference. This intervention reduced the average gating probability for the top-4 experts by 54% and led to a 29% drop in mean reciprocal rank (MRR), demonstrating that disabling a single key head disrupts downstream expert routing.
@@
In a complementary experiment, we forcibly activated expert $E^{20}_{59}$ regardless of the gating probabilities. This intervention recovered 85% of the original MRR, indicating that reinstating the correct expert can largely restore model performance even when upstream signals are perturbed.
@@
We applied Integrated Gradients to trace the decision path from attention heads to experts. On average, 28% of the attribution for expert activations (e.g. the route $H^{22}_{13}\!\to\!E^{21}_{48}$) could be directly linked back to specific attention heads, quantifying their direct influence in the routing computation.
@@
Table 4: Top 3 Attention Heads and Experts in Qwen 1.5-MoE Model with High Importance Scores and Routing Frequency.
5 Expert Specialization and Architectural Trade-offs
In this section, we investigate the phenomenon of neuron-level knowledge localization in MoE models. Our analysis focuses on the top 100 most important neurons, and the experts to whom they belong. Here, our analysis is more concentrated on fine-grained experts, whose specialization trends are particularly pronounced (Wang et al., 2024); for coarse-grained MoE models such as Mixtral, we provide a detailed description in Appendix A. We provide insights into expert specialization (§ 5.1) and investigate how architectural design affects the robustness and redundancy of the model, especially under expert failure scenarios (§ 5.2). Through systematic expert-blocking experiments, we reveal critical trade-offs between specialization depth, task sensitivity, and fault tolerance.
@@ 5 Expert Specialization and Architectural Trade-offs
Table 5: Top 5 most frequently selected experts for each knowledge type, along with the MRR of the model after excluding the top 1, 5, and 10 experts. The complete experimental results can be found in Appendix A.
@@ 5.2 Robustness and Redundancy in MoE Architectures
Table 5 shows the changes in MRR of Qwen 1.5-MoE and OLMoE models after blocking high-frequency experts under different knowledge types. Figure 7 visualizes the heat map of the MRR drop of the two models under different settings. Experimental results show that as the Top-1, Top-5, and Top-10 experts are blocked, Qwen 1.5-MoE and OLMoE show different robustness in different tasks.
@@ Results
Figure 5 shows the HIT@10 and MRR of different models on 12 tasks. It can be seen that the performance of different models varies greatly in some types. Therefore, when analyzing the experimental results, we mainly focus on the relations with similar performance to avoid the influence of the model’s internal parameter knowledge on the effect of knowledge attribution.
@@ Appendix B Detailed Analysis of Coarse/Fine-Grained Routed Experts
We report the full expert-isolation results in Table 8. As expected, the two fine-grained MoE models suffer precipitous drops in performance when their top specialists are blocked—reaffirming the strong localization of functionality we described in § 5.2. By contrast, Mixtral—a coarse-grained MoE with only eight experts per layer and a top-2 routing policy—maintains near-identical accuracy even when its ten most “important” experts are silenced.

### v2

@@ Abstract
The interpretability of Mixture-of-Experts (MoE) models, especially those with heterogeneous designs, remains underexplored. Existing attribution methods for dense models fail to capture dynamic routing-expert interactions in sparse MoE architectures. To address this issue, we propose a cross-level attribution algorithm to analyze sparse MoE architectures (Qwen 1.5-MoE, OLMoE, Mixtral-8x7B) against dense models (Qwen 1.5-7B, Llama-7B, Mistral-7B). Results show MoE models achieve 31% higher per-layer efficiency via a “mid-activation, late-amplification” pattern: early layers screen experts, while late layers refine knowledge collaboratively. Ablation studies reveal a “basic-refinement” framework—shared experts handle general tasks (entity recognition), while routed experts specialize in domain-specific processing (geographic attributes). Semantic-driven routing is evidenced by strong correlations between attention heads and experts ($r=0.68$), enabling task-aware coordination. Notably, architectural depth dictates robustness: deep Qwen 1.5-MoE mitigates expert failures (e.g., 43% MRR drop in geographic tasks when blocking top-10 experts) through shared expert redundancy, whereas shallow OLMoE suffers severe degradation (76% drop). Task sensitivity further guides design: core-sensitive tasks (geography) require concentrated expertise, while distributed-tolerant tasks (object attributes) leverage broader participation. These insights advance MoE interpretability, offering principles to balance efficiency, specialization, and robustness.
@@ 1 Introduction
Our main contributions are as follows:
@@ 2.1 MoE Architecture
The final hidden states $\bm{h}_{i}^{L}$ need to go through the unembedding process to map the high-dimensional representations back to the vocabulary space. Specifically, for each position $i$, the final output token vector $\bm{y}_{i}$ is obtained by multiplying the hidden state $\bm{h}_{i}^{L}$ with the unembedding matrix $\bm{W}_{\text{unembed}}$ and passing it through a softmax function.
@@ 2.2 Neuron-Level Attribution
We adopt the neuron-level attribution methods from Yu and Ananiadou (2024) for the FFN and attention layers and extend them for MoE models. For attention neurons $\bm{v}^{l}_{A}$, a vector in $l$-th Attention layer, the importance score is:
$\mathcal{I}(\bm{v}^{l}_{A})=\log p(x_{i}|\bm{v}^{l}_{A}+\bm{h}^{l-1})-\log p(x_{i}|\bm{h}^{l-1}),$ (7)
@@
where $\bm{v}^{l}_{F}$ is a vector (neuron) in $l$-th FFN layer.
@@
where $g_{i,j}^{l}$ is the gating probability for expert $\mathcal{E}_{j}$, and $v^{l}_{\mathcal{E}_{j}}$ is the neuron from expert $\mathcal{E}_{j}$ of $l$-th layer.
@@ 3.1 Layer Efficiency and Knowledge Concentration in MoE Models
Table 1 summarizes the performance of the four models, revealing that MoE architectures achieve notably higher per-layer gains. For example, Qwen 1.5-MoE attains a layer efficiency of 0.307 using only 24 layers, whereas its dense counterpart, Qwen 1.5-7B, requires 32 layers to reach an efficiency of just 0.203. Despite a slightly lower overall FFN gain, OLMoE—by virtue of its shallower design—achieves the highest layer efficiency at 0.311, demonstrating a marked optimization in computational density.
The data further indicate that MoE models concentrate their knowledge processing at earlier stages. Specifically, Qwen 1.5-MoE and OLMoE reach their peak FFN gains at 84.8% and 84.6% of the network depth, respectively. This “early routing, late refinement” pattern enables these models to perform initial expert filtering in mid-layers and then focus on intensive knowledge refinement in later layers. In contrast, dense models such as Qwen 1.5-7B and Llama-7B display a more gradual accumulation of gains—peaking at 90.2% and 77.6% of the network depth—which, while beneficial for cross-layer integration, tends to blur the functional distinctions among layers and reduce interpretability.
@@
Furthermore, a comparison of attention versus FFN gains underscores divergent processing strategies. Qwen 1.5-MoE, for instance, exhibits a pronounced reliance on specialized expert networks, with an FFN gain of 7.36 contrasted against an attention gain of only 3.18. Dense models, by comparison, distribute gains more evenly (with attention gains of 4.14 and 4.03 for Qwen 1.5-7B and Llama-7B, respectively), complicating the attribution of specific processing roles to individual layers. Notably, OLMoE achieves nearly identical attention and FFN gains (4.98 and 4.40, respectively), reflecting a balanced strategy that integrates both attention mechanisms and expert routing.
@@ 3.2 FFN Gain Evolution: Mid-Stage Activation and Late-Stage Amplification
Figure 3: Cumulative FFN layers contribution curves.
@@ 4.3 Causality Analysis
We selectively suppressed the activity of head $\mathcal{H}^{19}_{12}$ in Qwen during inference. This intervention reduced the average gating probability for the top-4 experts by 54% and led to a 29% drop in mean reciprocal rank (MRR), demonstrating that disabling a single key head disrupts downstream expert routing.
@@
In a complementary experiment, we forcibly activated expert $\mathcal{E}^{20}_{59}$ regardless of the gating probabilities. This intervention recovered 85% of the original MRR, indicating that reinstating the correct expert can largely restore model performance even when upstream signals are perturbed.
@@
We applied Integrated Gradients to trace the decision path from attention heads to experts. On average, 28% of the attribution for expert activations (e.g. the route $\mathcal{H}^{22}_{13}\to\mathcal{E}^{21}_{48}$) could be directly linked back to specific attention heads, quantifying their direct influence in the routing computation.
@@
Table 4: Top 3 attention heads and experts in Qwen 1.5-MoE.
5 MoE Knowledge Localization: Trade-offs and Robustness
In this section, we investigate the phenomenon of neuron-level knowledge localization in MoE models. For each knowledge type, identify the 100 most important neurons, determine which experts they belong to, and then we list the top 5 experts (i.e. the five experts containing the greatest number of those 100 neurons). Here, our analysis is more concentrated on fine-grained experts, whose specialization trends are particularly pronounced (Wang et al., 2024); for coarse-grained MoE models such as Mixtral, we provide a detailed description in Appendix B. We provide insights into expert specialization (§ 5.1) and investigate how architectural design affects the robustness and redundancy of the model, especially under expert failure scenarios (§ 5.2). Through systematic expert-blocking experiments, we reveal critical trade-offs between specialization depth, task sensitivity, and fault tolerance.
@@ 5 MoE Knowledge Localization: Trade-offs and Robustness
Table 5: Top 5 experts for each knowledge type, along with the model’s MRR after excluding the top 1, top 5, and top 10 of those experts. The complete experimental results can be found in Appendix B.
@@ 5.2 Robustness and Redundancy in MoE Architectures
Table 5 shows the changes in MRR of Qwen 1.5-MoE and OLMoE models after blocking high-frequency experts under different knowledge types. Experimental results show that as the Top-1, Top-5, and Top-10 experts are blocked, Qwen 1.5-MoE and OLMoE show different robustness in different tasks.
@@ Results
Figure 5 shows the HIT@10 of different models on 12 tasks. It can be seen that the performance of different models varies greatly in some types. Therefore, when analyzing the experimental results, we mainly focus on the relations with similar performance to avoid the influence of the model’s internal parameter knowledge on the effect of knowledge attribution.
@@ Appendix B Detailed Analysis of Coarse/Fine-Grained Routed Experts
We report the full expert-isolation results in Table 8. Figure 7 visualizes the heat map of the MRR drop of models under different settings. As expected, the two fine-grained MoE models suffer precipitous drops in performance when their top specialists are blocked—reaffirming the strong localization of functionality we described in § 5.2. By contrast, Mixtral—a coarse-grained MoE with only eight experts per layer and a top-2 routing policy—maintains near-identical accuracy even when its ten most “important” experts are silenced.